# D³-Predictor: Noise-Free Deterministic Diffusion for Dense Prediction

### v3

@@ 4.2.1 Depth Estimation
Depth estimation predicts the distance between the object and the camera from an image. We perform evaluations on five zero-shot real-world benchmarks, including NYUv2 [silberman2012indoor], KITTI [geiger2013vision], ETH3D [schops2017multi], ScanNet [dai2017scannet], and DIODE [vasiljevic1908diode]. The evaluation metrics include the Absolute Mean Relative Error (RelAbs) and the $\delta_{1}$-accuracy.
We compare specialized discriminative models [hu2024metric3d], multi-task models [le2025one], and diffusion-based models [ke2024repurposing]. Table 1 presents the quantitative comparison results. D³-Predictor demonstrates remarkable zero-shot generalization and promising performance across five benchmarks. Compared with diffusion-based methods, D³-Predictor requires neither multi-step sampling nor test-time ensembling, thereby achieving significantly higher inference efficiency and stability without sacrificing accuracy. Although our model lags behind state-of-the-art discriminative models in terms of quantitative performance, it produces visually sharper results, as shown in Fig. 4. When encountering out-of-domain scenes (e.g., AIGC-generated images), the performance of discriminative models drops drastically. Notably, most discriminative models require large-scale datasets and complex data pipelines, while our model only relies on limited synthetic data and arbitrary real images.
@@ 4.2.2 Surface Normal Estimation
Surface normal estimation predicts the direction vector perpendicular to the surface at each point on the object. We perform evaluations on five zero-shot real-world benchmarks, including NYUv2 [silberman2012indoor], ScanNet [dai2017scannet], iBims-1 [koch2018evaluation], DIODE [vasiljevic1908diode] and OASIS [chen2020oasis]. The evaluation metrics include mean angular error and the percentage of pixels with an angular error below the threshold of $11.25^{\circ}$.
We compare specialized discriminative models [hu2024metric3d] and diffusion-based models [ke2025marigold]. The quantitative results in Table 2 reveal that our D³-Predictor is competitive with state-of-the-art diffusion-based models across five benchmarks, while also demonstrating impressive inference efficiency and stability. Although E2E-FT [garcia2025fine] achieves marginally better performance than D³-Predictor, our model is significantly more data-efficient, requiring less than half of its training data. This is particularly advantageous in real-world applications, where annotating large-scale dense labels and training diffusion models is costly and thus often limits deployment. Fig. 4 further highlights the geometric detail generation capability and robustness of our approach.
@@ 4.2.3 Image Matting
Image matting aims to extract the foreground and background from an image and generate an alpha matte. We evaluate our method on the P3M-500-NP test set [li2021privacy]. The evaluation metrics include the sum of absolute differences (SAD), mean squared error (MSE), mean absolute difference (MAD) and connectivity (Conn.).
Table 3 presents the quantitative comparison results. Overall, our method significantly outperforms a recent diffusion-based method GenPercept [xu2024matters], demonstrating the superiority of our D³-Predictor for image matting. Moreover, compared with state-of-the-art discriminative matting models, D³-Predictor achieves clearly competitive performance. Qualitative comparisons in Fig. 5 further show that our method produces high-quality alpha mattes, especially in challenging regions such as transparent area and hair.
@@
HATT [qiao2020attention] 60K 30.35 0.0176 0.0072 27.42 4.8
SHM [chen2018semantic] 53K 20.77 0.0122 0.0093 17.09 4.3
MODNet [ke2022modnet] 45K 16.70 0.0097 0.0051 13.81 3
P3M-Net [li2021privacy] 9K 11.23 0.0065 0.0035 12.51 2
ViTAE-S [ma2023rethinking] 9K 7.59 0.0044 0.0019 6.96 1
GenPercept [xu2024matters] 9K 12.77 0.0074 0.0027 10.46 2
@@ 4.2.4 Generalization to Domain-Specific Tasks
To further assess the generalization and practical utility of our D³-Predictor, we evaluate it on 20 additional domain-specific dense prediction tasks, such as medical image analysis. These tasks are typically constrained by data scarcity, often due to factors like privacy protection. For these evaluations, we train our model using only $15$ samples per task with a simple MSE loss. We present Spinal Morphology Assessment source from [chu2015annotated] as a representative example.
As summarized in Table 4, D³-Predictor significantly outperforms strong zero-shot baselines, such as SAM [kirillov2023segment], by over $\mathbf{60\%}$ in average performance, demonstrating its superior generalization in the medical domain. The qualitative results in Fig. 6 further show our model accurately captures the complex structure of the spine, underscoring its excellent fine-grained geometric perception. More comprehensive results on the remaining tasks are in the Appendix.
@@
SAM [kirillov2023segment] 1.1B 0.555 0.611 0.612 0.593
CLIPSeg [luddecke2022image] 345K 0.519 0.569 0.551 0.546
Grounded-SAM [ren2024grounded] 1.1B 0.464 0.481 0.481 0.475
@@ 4.3 Ablation Study
For simplicity, we conduct ablation studies on the representative task of depth estimation. We report results on two standard benchmarks: NYUv2 [silberman2012indoor] for indoor scenes and KITTI [geiger2013vision] for driving scenes.
@@ 4.3.1 Training Data Volume
Our method achieves a remarkable balance between training data volume and performance. To further evaluate training data efficiency, we train D³-Predictor on varying amounts of synthetic data. As shown in Fig. 8, D³-Predictor already delivers acceptable performance even with a limited amount of training data, demonstrating the training data efficiency of our approach. It is worth noting that compared with state-of-the-art discriminative depth estimation models, such as Depth Pro [bochkovskii2024depth], our method exhibits superior data efficiency. First, Depth Pro is trained using a combination of more than twenty heterogeneous datasets, whereas our model only requires a small subset from three datasets. Second, Depth Pro relies on complex data augmentation pipelines to utilize large-scale mixed data, while our method does not require any specialized data processing.
@@ 4.3.2 Improving Generalization with Real Images
As illustrated in Fig. 8, further increasing the amount of synthetic training data to 75K yields only marginal performance gains, suggesting that the limited diversity of synthetic data constrains model generalization. To address this limitation, we train our model on a balanced mixture of synthetic data and real images with pseudo-labels to improve its generalization. This approach is feasible for two main reasons. First, state-of-the-art discriminative models, although less capable of producing fine-grained visual details than diffusion-based methods, typically achieve superior quantitative performance and can generate reliable pseudo-labels for real images. Second, diffusion priors, including our noise-free aggregated diffusion prior, inherently originate from Internet-scale datasets [schuhmann2022laion], which enables our model to effectively process both synthetic and real data in a unified manner. Table 5 demonstrates that incorporating real images significantly improves the model’s generalization ability across zero-shot real-world benchmarks.
@@ 4.3.3 Architecture of Diffusion Model
Recent progress in image generation has shifted from the U-Net architecture to the DiT architecture [peebles2023scalable], achieving better image quality and fidelity [flux2024, labs2025flux1kontextflowmatching, wu2025qwen]. We also initialize our D³-Predictor with FLUX.1 dev [flux2024], a DiT-based pretrained diffusion model. The detailed design of the DiT-based D³-Predictor is provided in the Appendix.
Due to the significant increase in model size (over $15\times$) and our limited computational resources, we perform only a quick validation on 7.5K synthetic samples using an input resolution of $\times768$. Note that this resolution is sub-optimal for FLUX.1 dev [flux2024], which is pretrained on much higher resolutions. As shown in Table 6, the DiT-based D³-Predictor achieves competitive quantitative performance compared to the U-Net variant. Furthermore, Fig. 7 reveals that it exhibits higher perceptual fidelity. These findings suggest that the DiT-based D³-Predictor holds immense potential for advancing dense prediction performance.
@@ 5 Conclusion
Note: An anonymous static site with our visualization results is available at ./index.html. We will keep updating it and release it publicly with our resources.
@@ 6.1.1 U-Net-Based D³-Predictor
The U-Net-based D³-Predictor is initialized using the pretrained Stable Diffusion v2.1 (SD v2.1) [rombach2022high]. Figure 9 illustrates the backbone architecture of SD v2.1 and highlights the specific internal representations selected for self-supervised alignment (indicated by dashed arrows).
The selection of layers for representation alignment is critical. Prior work [tumanyan2023plug] suggests that SD’s upsampling blocks encode rich spatial semantic information that is largely invariant to appearance features (e.g., style and texture), focusing instead on underlying structure and shape. Inspired by this finding, we hypothesize that the internal representations from SD’s upsampling blocks are optimally suited for dense prediction, given that this task is fundamentally an image-to-structure mapping [fu2024geowizard]. To maintain efficiency without compromising alignment quality, we carefully chose a minimal set of layers. This selection captures sufficient multi-scale structural detail while avoiding the high computational cost of aligning redundant representations. Consequently, we select the specific internal representations indicated by the dashed arrows in Figure 9.
@@ 6.1.2 DiT-Based D³-Predictor
The DiT-based D³-Predictor is initialized using the pretrained FLUX.1 dev model [flux2024]. Figure 10 presents the backbone architecture of FLUX.1 dev and indicates the internal representations selected for the self-supervised alignment (marked by dashed arrows). Drawing upon the insights from the U-Net-based D³-Predictor implementation, we focus on the representations from the deeper layers which typically carry refined spatial and structural information in DiT architectures. Specifically, we select the representations from the last $25$ single-flow blocks for robust and effective representation alignment.
@@ 6.2 Text Prompt
Text prompts provide diverse guidance for creative image [li2024photomaker, xu2024diffusion, qin2023unicontrol] and video [yang2024cogvideox, zhao2024moviedreamer, wang2024framer] generation, ensuring the results align with user intent. Despite their utility in creative tasks, extensive prior works [ke2024repurposing, xu2024matters, zhao2025diception, ke2025marigold] suggest that complex text prompts are unnecessary for diffusion-based dense prediction models. This is primarily because dense prediction is a deterministic mapping from image to geometric structure [fu2024geowizard], which does not require textual guidance for diversity. Consequently, the D³-Predictor utilizes a simple, fixed text prompt for each task to efficiently maintain the necessary context:
@@ 6.3 Training Datasets Mixing Strategy
As described in Section $4.1.1$ of our main paper, we utilize multiple training datasets for both depth and surface normal estimation. Inspired by previous work [garcia2025fine, fu2024geowizard], we implement the following dataset sampling configuration in each iteration: For depth estimation, we sample Hypersim [roberts2021hypersim] at a rate of 0.76, Virtual KITTI [cabon2020virtual] at 0.12, and COCO [lin2014microsoft] with pseudo-depth labels generated by Depth Pro [bochkovskii2024depth] at 0.12; for surface normal estimation, we sample Hypersim [roberts2021hypersim] at a rate of 0.9 and Sintel [butler2012naturalistic] at 0.1.
@@ 6.4 Training Objective
The first part of the objective is the Self-Supervised Diffusion Priors Aggregation, which is implemented via the alignment of internal representations. Given that prior works [yu2024representation, wu2025representation] have demonstrated the established efficacy of negative cosine similarity for representation alignment in diffusion models, we adopt it as the alignment metric $\text{dist}(\cdot,\cdot)$ in Equation $2$.
@@
For depth estimation, $\mathcal{L}_{\text{task}}$ is formulated as a weighted combination of three components: the MSE loss $\mathcal{L}_{\text{MSE}}$, the affine-invariant loss $\mathcal{L}_{\text{aff}}$~[ranftl2020towards], and the gradient loss in log space $\mathcal{L}_{\text{grad}}$~[ranftl2020towards]. Specifically, $\mathcal{L}_{\text{task}}$ is defined as:
@@ 8.1 Ablation on Projection Head
We conduct an ablation study to investigate the contribution of the timestep-conditioned projection head used to map the internal representations of the D³-Predictor. We evaluate the performance of the D³-Predictor when trained with and without the projection head. Following the setup detailed in Section $4.3$ of our main paper, we report results on two standard depth estimation benchmarks: NYUv2 [silberman2012indoor] for indoor scenes and KITTI [geiger2013vision] for driving scenes. The comparison results, shown in Table 7, consistently demonstrate that the D³-Predictor trained with the projection head outperforms the model trained without it on both benchmarks. This indicates that the projection head is essential for effectively bridging the representation of the D³-Predictor and the timestep-dependent representations extracted from the pretrained diffusion model.
@@ 8.3 Efficiency and Runtime Analysis
We conduct a thorough analysis of the efficiency of the D³-Predictor, summarizing the results in Table 9. The reported inference time is the average runtime over $5,000$ images at $768\times 768$ resolution on a single NVIDIA L40S GPU. Our D³-Predictor achieves the fastest inference speed among all evaluated baselines. Specifically, our method is approximately $66\%$ faster than the discriminative model Metric3Dv2 [hu2024metric3d] and achieves nearly a $10\%$ speedup over the fastest comparable diffusion-based methods, such as E2E-FT [garcia2025fine] and Lotus-G [he2024lotus].
Beyond speed, the D³-Predictor demonstrates highly competitive GPU memory efficiency. While Lotus-G [he2024lotus] maintains the minimum memory consumption, our model’s usage is comparable and represents a significant saving of nearly $50\%$ compared to the most memory-intensive baseline, Depth Pro [bochkovskii2024depth]. Even when compared to other efficient diffusion-based models like E2E-FT [garcia2025fine], we still achieve over $20\%$ savings in memory consumption. This excellent performance in both inference time and memory requirement confirms that the D³-Predictor is an efficiency diffusion-based method while maintaining high predictive performance.
@@
Metric3Dv2 [hu2024metric3d] $\sim 0.66$ $7.5$
Depth Pro [bochkovskii2024depth] $\sim 0.64$ $12.2$
@@
E2E-FT [garcia2025fine] $\sim 0.27$ $8.07$
Marigold-Normals [ke2025marigold] $\sim 0.45$ $8.04$
Lotus-G [he2024lotus] $\sim 0.25$ $5.72$
GenPercept [xu2024matters] $\sim 0.28$ $8.05$

### v4

@@ 4.2.1 Depth Estimation
Depth estimation predicts the distance between the object and the camera from an image. We perform evaluations on five zero-shot real-world benchmarks, including NYUv2 [69], KITTI [19], ETH3D [67], ScanNet [13], and DIODE [73]. The evaluation metrics include the Absolute Mean Relative Error (RelAbs) and the $\delta_{1}$-accuracy.
We compare specialized discriminative models [25], multi-task models [40], and diffusion-based models [31]. Table 1 presents the quantitative comparison results. D³-Predictor demonstrates remarkable zero-shot generalization and promising performance across five benchmarks. Compared with diffusion-based methods, D³-Predictor requires neither multi-step sampling nor test-time ensembling, thereby achieving significantly higher inference efficiency and stability without sacrificing accuracy. Although our model lags behind state-of-the-art discriminative models in terms of quantitative performance, it produces visually sharper results, as shown in Fig. 4. When encountering out-of-domain scenes (e.g., AIGC-generated images), the performance of discriminative models drops drastically. Notably, most discriminative models require large-scale datasets and complex data pipelines, while our model only relies on limited synthetic data and arbitrary real images.
@@ 4.2.2 Surface Normal Estimation
Surface normal estimation predicts the direction vector perpendicular to the surface at each point on the object. We perform evaluations on five zero-shot real-world benchmarks, including NYUv2 [69], ScanNet [13], iBims-1 [36], DIODE [73] and OASIS [8]. The evaluation metrics include mean angular error and the percentage of pixels with an angular error below the threshold of $11.25^{\circ}$.
We compare specialized discriminative models [25] and diffusion-based models [32]. The quantitative results in Table 2 reveal that our D³-Predictor is competitive with state-of-the-art diffusion-based models across five benchmarks, while also demonstrating impressive inference efficiency and stability. Although E2E-FT [18] achieves marginally better performance than D³-Predictor, our model is significantly more data-efficient, requiring less than half of its training data. This is particularly advantageous in real-world applications, where annotating large-scale dense labels and training diffusion models is costly and thus often limits deployment. Fig. 4 further highlights the geometric detail generation capability and robustness of our approach.
@@ 4.2.3 Image Matting
Image matting aims to extract the foreground and background from an image and generate an alpha matte. We evaluate our method on the P3M-500-NP test set [43]. The evaluation metrics include the sum of absolute differences (SAD), mean squared error (MSE), mean absolute difference (MAD) and connectivity (Conn.).
Table 3 presents the quantitative comparison results. Overall, our method significantly outperforms a recent diffusion-based method GenPercept [80], demonstrating the superiority of our D³-Predictor for image matting. Moreover, compared with state-of-the-art discriminative matting models, D³-Predictor achieves clearly competitive performance. Qualitative comparisons in Fig. 5 further show that our method produces high-quality alpha mattes, especially in challenging regions such as transparent area and hair.
@@
HATT [57] 60K 30.35 0.0176 0.0072 27.42 4.8
SHM [7] 53K 20.77 0.0122 0.0093 17.09 4.3
MODNet [33] 45K 16.70 0.0097 0.0051 13.81 3
P3M-Net [43] 9K 11.23 0.0065 0.0035 12.51 2
ViTAE-S [49] 9K 7.59 0.0044 0.0019 6.96 1
GenPercept [80] 9K 12.77 0.0074 0.0027 10.46 2
@@ 4.2.4 Generalization to Domain-Specific Tasks
To further assess the generalization and practical utility of our D³-Predictor, we evaluate it on 20 additional domain-specific dense prediction tasks, such as medical image analysis. These tasks are typically constrained by data scarcity, often due to factors like privacy protection. For these evaluations, we train our model using only $15$ samples per task with a simple MSE loss. We present Spinal Morphology Assessment source from [12] as a representative example.
As summarized in Table 4, D³-Predictor significantly outperforms strong zero-shot baselines, such as SAM [35], by over $\mathbf{60\%}$ in average performance, demonstrating its superior generalization in the medical domain. The qualitative results in Fig. 6 further show our model accurately captures the complex structure of the spine, underscoring its excellent fine-grained geometric perception. More comprehensive results on the remaining tasks are in the Appendix.
@@
SAM [35] 1.1B 0.555 0.611 0.612 0.593
CLIPSeg [48] 345K 0.519 0.569 0.551 0.546
Grounded-SAM [62] 1.1B 0.464 0.481 0.481 0.475
@@ 4.3 Ablation Study
For simplicity, we conduct ablation studies on the representative task of depth estimation. We report results on two standard benchmarks: NYUv2 [69] for indoor scenes and KITTI [19] for driving scenes.
@@ 4.3.1 Training Data Volume
Our method achieves a remarkable balance between training data volume and performance. To further evaluate training data efficiency, we train D³-Predictor on varying amounts of synthetic data. As shown in Fig. 8, D³-Predictor already delivers acceptable performance even with a limited amount of training data, demonstrating the training data efficiency of our approach. It is worth noting that compared with state-of-the-art discriminative depth estimation models, such as Depth Pro [4], our method exhibits superior data efficiency. First, Depth Pro is trained using a combination of more than twenty heterogeneous datasets, whereas our model only requires a small subset from three datasets. Second, Depth Pro relies on complex data augmentation pipelines to utilize large-scale mixed data, while our method does not require any specialized data processing.
@@ 4.3.2 Improving Generalization with Real Images
As illustrated in Fig. 8, further increasing the amount of synthetic training data to 75K yields only marginal performance gains, suggesting that the limited diversity of synthetic data constrains model generalization. To address this limitation, we train our model on a balanced mixture of synthetic data and real images with pseudo-labels to improve its generalization. This approach is feasible for two main reasons. First, state-of-the-art discriminative models, although less capable of producing fine-grained visual details than diffusion-based methods, typically achieve superior quantitative performance and can generate reliable pseudo-labels for real images. Second, diffusion priors, including our noise-free aggregated diffusion prior, inherently originate from Internet-scale datasets [68], which enables our model to effectively process both synthetic and real data in a unified manner. Table 5 demonstrates that incorporating real images significantly improves the model’s generalization ability across zero-shot real-world benchmarks.
@@ 4.3.3 Architecture of Diffusion Model
Recent progress in image generation has shifted from the U-Net architecture to the DiT architecture [56], achieving better image quality and fidelity [38, 37, 77]. We also initialize our D³-Predictor with FLUX.1 dev [38], a DiT-based pretrained diffusion model. The detailed design of the DiT-based D³-Predictor is provided in the Appendix.
Due to the significant increase in model size (over $15\times$) and our limited computational resources, we perform only a quick validation on 7.5K synthetic samples using an input resolution of $\times768$. Note that this resolution is sub-optimal for FLUX.1 dev [38], which is pretrained on much higher resolutions. As shown in Table 6, the DiT-based D³-Predictor achieves competitive quantitative performance compared to the U-Net variant. Furthermore, Fig. 7 reveals that it exhibits higher perceptual fidelity. These findings suggest that the DiT-based D³-Predictor holds immense potential for advancing dense prediction performance.
@@ 6.1.1 U-Net-Based D³-Predictor
The U-Net-based D³-Predictor is initialized using the pretrained Stable Diffusion v2.1 (SD v2.1) [66]. Figure 9 illustrates the backbone architecture of SD v2.1 and highlights the specific internal representations selected for self-supervised alignment (indicated by dashed arrows).
The selection of layers for representation alignment is critical. Prior work [72] suggests that SD’s upsampling blocks encode rich spatial semantic information that is largely invariant to appearance features (e.g., style and texture), focusing instead on underlying structure and shape. Inspired by this finding, we hypothesize that the internal representations from SD’s upsampling blocks are optimally suited for dense prediction, given that this task is fundamentally an image-to-structure mapping [17]. To maintain efficiency without compromising alignment quality, we carefully chose a minimal set of layers. This selection captures sufficient multi-scale structural detail while avoiding the high computational cost of aligning redundant representations. Consequently, we select the specific internal representations indicated by the dashed arrows in Figure 9.
@@ 6.1.2 DiT-Based D³-Predictor
The DiT-based D³-Predictor is initialized using the pretrained FLUX.1 dev model [38]. Figure 10 presents the backbone architecture of FLUX.1 dev and indicates the internal representations selected for the self-supervised alignment (marked by dashed arrows). Drawing upon the insights from the U-Net-based D³-Predictor implementation, we focus on the representations from the deeper layers which typically carry refined spatial and structural information in DiT architectures. Specifically, we select the representations from the last $25$ single-flow blocks for robust and effective representation alignment.
@@ 6.2 Text Prompt
Text prompts provide diverse guidance for creative image [44, 79, 58] and video [83, 91, 75] generation, ensuring the results align with user intent. Despite their utility in creative tasks, extensive prior works [31, 80, 92, 32] suggest that complex text prompts are unnecessary for diffusion-based dense prediction models. This is primarily because dense prediction is a deterministic mapping from image to geometric structure [17], which does not require textual guidance for diversity. Consequently, the D³-Predictor utilizes a simple, fixed text prompt for each task to efficiently maintain the necessary context:
@@ 6.3 Training Datasets Mixing Strategy
As described in Section $4.1.1$ of our main paper, we utilize multiple training datasets for both depth and surface normal estimation. Inspired by previous work [18, 17], we implement the following dataset sampling configuration in each iteration: For depth estimation, we sample Hypersim [65] at a rate of 0.76, Virtual KITTI [6] at 0.12, and COCO [45] with pseudo-depth labels generated by Depth Pro [4] at 0.12; for surface normal estimation, we sample Hypersim [65] at a rate of 0.9 and Sintel [5] at 0.1.
@@ 6.4 Training Objective
The first part of the objective is the Self-Supervised Diffusion Priors Aggregation, which is implemented via the alignment of internal representations. Given that prior works [86, 78] have demonstrated the established efficacy of negative cosine similarity for representation alignment in diffusion models, we adopt it as the alignment metric $\text{dist}(\cdot,\cdot)$ in Equation $2$.
@@
For depth estimation, $\mathcal{L}_{\text{task}}$ is formulated as a weighted combination of three components: the MSE loss $\mathcal{L}_{\text{MSE}}$, the affine-invariant loss $\mathcal{L}_{\text{aff}}$~[61], and the gradient loss in log space $\mathcal{L}_{\text{grad}}$~[61]. Specifically, $\mathcal{L}_{\text{task}}$ is defined as:
@@ 8.1 Ablation on Projection Head
We conduct an ablation study to investigate the contribution of the timestep-conditioned projection head used to map the internal representations of the D³-Predictor. We evaluate the performance of the D³-Predictor when trained with and without the projection head. Following the setup detailed in Section $4.3$ of our main paper, we report results on two standard depth estimation benchmarks: NYUv2 [69] for indoor scenes and KITTI [19] for driving scenes. The comparison results, shown in Table 7, consistently demonstrate that the D³-Predictor trained with the projection head outperforms the model trained without it on both benchmarks. This indicates that the projection head is essential for effectively bridging the representation of the D³-Predictor and the timestep-dependent representations extracted from the pretrained diffusion model.
@@ 8.3 Efficiency and Runtime Analysis
We conduct a thorough analysis of the efficiency of the D³-Predictor, summarizing the results in Table 9. The reported inference time is the average runtime over $5,000$ images at $768\times 768$ resolution on a single NVIDIA L40S GPU. Our D³-Predictor achieves the fastest inference speed among all evaluated baselines. Specifically, our method is approximately $66\%$ faster than the discriminative model Metric3Dv2 [25] and achieves nearly a $10\%$ speedup over the fastest comparable diffusion-based methods, such as E2E-FT [18] and Lotus-G [23].
Beyond speed, the D³-Predictor demonstrates highly competitive GPU memory efficiency. While Lotus-G [23] maintains the minimum memory consumption, our model’s usage is comparable and represents a significant saving of nearly $50\%$ compared to the most memory-intensive baseline, Depth Pro [4]. Even when compared to other efficient diffusion-based models like E2E-FT [18], we still achieve over $20\%$ savings in memory consumption. This excellent performance in both inference time and memory requirement confirms that the D³-Predictor is an efficiency diffusion-based method while maintaining high predictive performance.
@@
Metric3Dv2 [25] $\sim 0.66$ $7.5$
Depth Pro [4] $\sim 0.64$ $12.2$
@@
E2E-FT [18] $\sim 0.27$ $8.07$
Marigold-Normals [32] $\sim 0.45$ $8.04$
Lotus-G [23] $\sim 0.25$ $5.72$
GenPercept [80] $\sim 0.28$ $8.05$